# Majorana fermion induced power-law scaling in the violations of the Wiedemann-Franz law

### v2

@@ IV.1 Inelastic scattering with both phase and momentum relaxation
In the first model (see Fig. 1 (a)), we use an inelastic scatterer with both phase and momentum relaxation [8]. The S-matrix for the inelastic scatterer with both phase and momentum relaxation is given by [8],
@@
the matrix for the STIM junction is given by Eq. (3a). Buttiker introduced the voltage probe model, as discussed in Ref. [8] to implement inelastic scattering with both phase and momentum relaxation and in Ref. [9] to implement inelastic scattering with only phase relaxation. In this model, the charge current vanishes at the voltage probe. Electrons lose their phase memory upon entering the voltage probe and are re-injected back into the sample with the same probability. Later, M. Kilgour and D. Segal extended this concept by introducing the voltage-temperature probe, as detailed in Refs. [10, 11, 12]. This generalization of the Buttiker voltage probe to thermoelectric transport ensures that both charge and heat currents vanish at the voltage-temperature probe. Experimentally, a voltage temerature probe can be implemented by first ensuring a purely coherent system. Additionally, the system must be tuned to the linear response regime (i.e., near equilibrium), which involves applying small bias voltages (on the order of millivolts) and small temperature biases (around 1 Kelvin). Experimentally, one can vary the inelastic scattering parameter $\epsilon$ via a gate voltage in a voltage probe as has been shown in Ref. [26].
@@
Using Landauer-Buttiker scattering theory, we can find the charge and heat currents in each terminal of the setup described in Fig. 1 (a). In our work, the left and right terminals are at voltages $V_{1}=0$, and $V_{2}=V$, respectively. Similarly, the temperatures of the left and right terminals are $T_{1}$ and $T_{1}+\Delta T_{R}$ respectively. The Buttiker voltage-temperature probe is at temperature $T+\Delta T$, and voltage $V_{3}$. From Eq. (11), we can write the charge and heat currents for our setup in terms of the Onsager elements as,
@@
We are solving Eqs. (3-8) and using Eqs. (12) for the S-matrix of the Buttiker voltage-temperature probe, one can find the transmission probabilities $\mathcal{T}^{ss;kk}_{ij}$ and $\mathcal{T}^{rs;lk}_{ij}$ in the setup. Plugging the transmission probabilities in Eqs. (13-16), along with Eq. (9), allows us to calculate the charge and heat currents in each terminal. We represent the charge and heat currents in terms of the voltage and temperature biases in a matrix notation below: The charge currents in each terminal due to spin-up electrons and holes are given by,
@@
where $G(X)$ is the electrical conductance for particles with transmission probability $X$, $S(X)$ is the Seebeck coefficient for particles with transmission probability $X$, and $L(X)$ is the thermal conductance for particles with transmission probability $X$. We can calculate the charge and heat currents for both the Buttiker S-matrix given in Eq. (12) by using the respective S-matrix for the Buttiker voltage-temperature probe [10] when solving for the transmission probabilities $\mathcal{T}^{ss;kk}_{ij}$ and $\mathcal{T}^{rs;lk}_{ij}$. In Eqs. (17-20) the electrical conductance $G(\mathcal{T}^{s;k}_{ij}(E))$ is given by,
@@
The total charge current in the Buttiker voltage-temperature probe is thus given by,
and the total heat current is given by,
In order to introduce inelastic scattering via the Buttiker voltage-temperature probe, we should have $I_{3c}(V,T+\Delta T)=I_{3q}(V,T+\Delta T)=0$. In our setup, we measure the charge and heat currents in the second terminal. The total charge current in the second terminal is given by,
and the total heat current is given by,
In order to find the thermoelectric coefficients and, subsequently, the Lorenz ratio, we write $I_{2q}$ and $I_{2c}$ in terms of $V$ and $\Delta T$ by eliminating $V_{3}$ and $\Delta T$ from Eqs. (17-20). We use the condition of the Buttiker voltage-temperature probe ($I_{3c}(V,T+\Delta T)=I_{3q}(V,T+\Delta T)=0$) in order to write all the voltages and temperatures in terms of $V$ and $\Delta T$. Thus, $I_{2c}$ and $I_{2q}$ can now be written solely in terms of $V$ and $\Delta T$ as,
$\small I_{2c}=L^{\prime}_{cV}V+L^{\prime}_{cT}\Delta T_{R}\text{ and }I_{2q}=L^{\prime}_{qV}V+L^{\prime}_{qT}\Delta T_{R},$ (28)
where $L^{\prime}_{cV},L^{\prime}_{qV},L^{\prime}_{cT},L^{\prime}_{qT}$ are the effective Onsager coefficients for conduction in the second terminal and are functions of $G(\mathcal{T}^{s;k}_{ij}(E))$, $S(\mathcal{T}^{s;k}_{ij}(E))$, $L(\mathcal{T}^{s;k}_{ij}(E))$ as defined in Eqs. (21-23). $L^{\prime}_{cV}$ is the effective electrical conductance in the second terminal given as, $L^{\prime}_{cV}=\frac{I_{2c}}{V}|_{\Delta T=0}.$ $L^{\prime}_{cT}$ is the electrical response to the temperature difference and is given by $L^{\prime}_{cT}=\frac{I_{2c}}{\Delta T_{R}}|_{V=0}$. Similarly, the thermal response to the voltage difference $L^{\prime}_{qV}$ is given by $L^{\prime}_{qV}=\frac{I_{2q}}{V}|_{\Delta T_{R}=0}$, and the thermal response to the temperature difference is given by $L^{\prime}_{qT}=\frac{I_{2q}}{\Delta T}|_{V=0}$. The effective Onsager coefficients are then given as,
where $\delta=\dfrac{1}{\sum_{s,k}\left[G_{33}^{s,k}L_{33}^{s,k}-(S_{33}^{s,k})^{2}\right]}$. The effective electrical conductance is the total charge current generated due to the voltage difference and is given by $\sigma^{\prime}=L^{\prime}_{cV}$. The effective thermal conductance [27] is the heat current generated by a unit temperature bias without any charge current. The effective thermal conductance of the setup with $3\times 3$ Buttiker probe is then given by,
$W=\frac{\kappa^{\prime}}{\sigma^{\prime}}.$ (34)
When WF law is preserved, $W=W_{0}$, where $W_{0}=\frac{\kappa_{0}}{G_{0}}$, where $G_{0}=\dfrac{e^{2}}{\hbar}$ and $\kappa_{0}=\pi^{2}k_{B}^{2}T/3h$, where $T$ is temperature of the reservoirs 1 and 2. WF law is derived from the fact that in condensed matter systems, both charge and heat are carried by the quasiparticles in the conductor, namely, the electrons and holes. Quasiparticles in conductors generally follow particle-hole symmetry (PHS), i.e., the electron energy levels are symmetric to the hole energy levels. When PHS is preserved in the system, WF law is preserved, and the Lorenz ratio is not violated [28, 29]. The breakdown of PHS causes violations of WF law. In the next subsection, we include inelastic scattering with the help of the Buttiker voltage-temperature probe but with phase relaxation only.
IV.2 Inelastic scattering with only phase relaxation
Fig. 1 (b) shows the MBS ABI with only phase relaxation [9]. The third terminal, i.e., the Buttiker voltage-temperature probe, is divided into two terminals that are connected to reservoirs at voltage $V_{3}$ and temperature $T+\Delta T$ such that the total charge current and heat current flowing in each terminal is zero. The inelastic scatterers are connected via a coupler described by a $4\times 4$ scattering matrix given by [9],

### v4

@@ IV.1 Inelastic scattering with both phase and momentum relaxation
In the first model (see Fig. 1 (a)), we use an inelastic scatterer with both phase and momentum relaxation [18]. The S-matrix for the inelastic scatterer with both phase and momentum relaxation is given below,
@@
where $p=\frac{1}{2}\left(\sqrt{1-2\epsilon}-1\right)$ and $p=\frac{1}{2}\left(\sqrt{1-2\epsilon}+1\right)$, see also Ref. [18]. When $\epsilon=0$, then complete backscattering occurs and the Aharanov-Bohm ring is disconnected from the Buttiker probe. It means no electron has entered into the Buttiker probe from Aharanov-Bohm ring, which indicates zero inelastic scattering. The BVT probe is a phenomenological method used to incorporate the effects of inelastic scattering in mesoscopic samples. Its primary role is to provide a theoretical foundation, using the single-electron transport-based Landauer-Buttiker model, to explain experimental results where electron-electron and electron-phonon scattering occur. This approach addresses the question of whether one can gain insight into how inelastic effects influence electron transport without performing an exact many-body calculation. This was the motivation behind Buttiker’s introduction of the voltage probe model, as discussed in Refs. [18, 19]. In this model, it is ensured that the charge current vanishes at the voltage probe. Electrons lose their phase memory upon entering the voltage probe and are re-injected back into the sample with the same probability. Later, M. Kilgour and D. Segal extended this concept by introducing the BVT probe, as detailed in Refs. [20, 21, 22]. This generalization of the Buttiker voltage probe to thermoelectric transport ensures that both charge and heat currents vanish at the BVT probe. Experimentally, a voltage probe can be implemented by ensuring a purely coherent system. Additionally, the system must be tuned to the linear response regime (i.e., near equilibrium), which involves applying small bias voltages (on the order of millivolts) and small temperature biases (around 1 Kelvin). Experimentally, one can also vary the inelastic scattering parameter $\epsilon$ in a BVT as has been performed in [34].
We can calculate the charge and heat currents for the S-matrix given in Eq. (12) by using the respective S-matrix for the BVT probe [20] after solving for the transmission probabilities $\mathcal{T}^{ss;kk}_{ij}$ and $\mathcal{T}^{rs;lk}_{ij}$, see, Appendix A. The total charge current in the BVT probe is thus given by,
and the total heat current is given by,
In order to introduce inelastic scattering via the BVT probe, we should have $I_{3c}(V,T+\Delta T)=I_{3q}(V,T+\Delta T)=0$. In our setup, we measure the charge and heat currents in the second terminal. The total charge current in the second terminal is given by,
and the total heat current is given by,
In order to find the thermoelectric coefficients and, subsequently, the Lorenz ratio, we write $I_{2q}$ and $I_{2c}$ in terms of $V$ and $\Delta T$ by eliminating $V_{3}$ and $\Delta T$ from Eqs. (36-39). We use the condition of the BVT probe ($I_{3c}(V,T+\Delta T)=I_{3q}(V,T+\Delta T)=0$) in order to write all the voltages and temperatures in terms of $V$ and $\Delta T$. Thus, $I_{2c}$ and $I_{2q}$ can now be written solely in terms of $V$ and $\Delta T$ as,
$I_{2c}=L^{\prime}_{cV}V+L^{\prime}_{cT}\Delta T_{R}\text{ and }I_{2q}=L^{\prime}_{qV}V+L^{\prime}_{qT}\Delta T_{R},$ (17)
where $L^{\prime}_{cV},L^{\prime}_{qV},L^{\prime}_{cT},L^{\prime}_{qT}$ are the effective Onsager coefficients for conduction in the second terminal and are functions of $G(\mathcal{T}^{s;k}_{ij}(E))$, $S(\mathcal{T}^{s;k}_{ij}(E))$, $L(\mathcal{T}^{s;k}_{ij}(E))$ as defined in Eqs. (40-42). $L^{\prime}_{cV}$ is the effective electrical conductance in the second terminal given as, $L^{\prime}_{cV}=\frac{I_{2c}}{V}|_{\Delta T=0}.$ $L^{\prime}_{cT}$ is the electrical response to the temperature difference and is given by $L^{\prime}_{cT}=\frac{I_{2c}}{\Delta T_{R}}|_{V=0}$. Similarly, the thermal response to the voltage difference $L^{\prime}_{qV}$ is given by $L^{\prime}_{qV}=\frac{I_{2q}}{V}|_{\Delta T_{R}=0}$, and the thermal response to the temperature difference is given by $L^{\prime}_{qT}=\frac{I_{2q}}{\Delta T}|_{V=0}$. The effective Onsager coefficients are then given as,
where $\delta=\dfrac{1}{\sum_{s,k}\left[G_{33}^{s,k}L_{33}^{s,k}-(S_{33}^{s,k})^{2}\right]}$. The effective electrical conductance is the total charge current generated due to the voltage difference and is given by $\sigma^{\prime}=L^{\prime}_{cV}$. The effective thermal conductance [35] is the heat current generated by a unit temperature bias without any charge current. The effective thermal conductance of the setup with Eq. (12) is then given by,
$W=\frac{\kappa^{\prime}}{\sigma^{\prime}}.$ (23)
When WF law is preserved, $W=W_{0}$, where $W_{0}=\frac{\kappa_{0}}{G_{0}}$, where $G_{0}=\dfrac{e^{2}}{\hbar}$ and $\kappa_{0}=\pi^{2}k_{B}^{2}T/3h$, where $T$ is temperature of the reservoirs 1 and 2. WF law is derived from the fact that in condensed matter systems, both charge and heat are carried by the quasiparticles in the conductor, namely, the electrons and holes. Quasiparticles in conductors generally follow particle-hole symmetry (PHS), i.e., the electron energy levels are symmetric to the hole energy levels. When PHS is preserved in the system, WF law is preserved, and the Lorenz ratio is not violated [36, 37]. The breakdown of PHS causes violations of WF law. In the next subsection, we include inelastic scattering with the help of the BVT probe but with phase relaxation only.
IV.2 Inelastic scattering with only phase relaxation
Fig. 1 (b) shows the MBS ABI with only phase relaxation [19]. The third terminal, i.e., the BVT probe, is divided into two terminals that are connected to reservoirs at voltage $V_{3}$ and temperature $T+\Delta T$ such that the total charge current and heat current flowing in each terminal is zero. The inelastic scatterers are connected via a coupler described by a $4\times 4$ scattering matrix given by [19],
where $\epsilon\in[0,1]$, and $I$ is the $2\times 2$ scattering matrix. $\epsilon=1$ denotes maximal coupling, while $\epsilon\to 0$ denotes no coupling [19]. Similar to the derivation of the Onsager matrix elements for the inelastic scatterer with both phase and momentum relaxation, one can find the Onsager matrix elements for the inelastic scatterer with only phase relaxation, see Appendix B. First, we write charge and heat current in Eqs. (47-50) in terms of $V$, and $\Delta T_{R}$ only by using the condition of the BVT probe ($I_{3c}(V,T+\Delta T)=I_{3q}(V,T+\Delta T)=I_{4c}(V,T+\Delta T)=I_{4q}(V,T+\Delta T)=0$). This allows us to write $I_{2c}$, and $I_{2q}$ in terms of $V$ and $\Delta T$ only as,
$\begin{split}I_{2c}=L^{\prime\prime}_{cV}V+L^{\prime\prime}_{cT}\Delta T_{R}\text{, and }I_{2q}=L^{\prime\prime}_{qV}V+L^{\prime\prime}_{qT}\Delta T_{R}.\\ \end{split}$ (25)
Figure 2: Lorentz ratio ($W$) vs. the coupling strength ($\epsilon$) of the BVT probe for (a) with both phase and momentum relaxation (b) with phase relaxation. Parameters are $E_{F}=10\mu eV$, $E_{M}=0.3\mu eV$, $\Gamma=1\mu eV$, $\phi=\phi_{0}$ and elastic scattering at junction $J_{1},J_{2}$, where $\phi=hc/e$ and $\alpha=\frac{1}{2}$ and represents elastic scattering at J1 and J2.
where, $\Delta=\sum_{s,k}\bigg{[}(G_{33}^{s,k}+G_{34}^{s,k}+G_{43}^{s,k}+G_{44}^{s,k})(L_{33}^{s,k}+L_{34}^{s,k}+L_{43}^{s,k}+L_{44}^{s,k})-(S_{33}^{s,k}+S_{34}^{s,k}+S_{43}^{s,k}+S_{44}^{s,k})^{2}\bigg{]}$.
The effective Onsager matrix elements can find all the thermoelectric coefficients. For the setup with the BVT probe with phase relaxation only, the thermal conductance is given by [38, 39],
The Lorenz ratio for the setup with the BVT probe with phase relaxation only is then given by,
The Mathematica codes are available in GitHub  [40].
V Results
Figure 3: Effect of change in elastic scattering: Lorentz ratio ($W$) vs. the coupling strength ($\epsilon$) of the BVT probe for (a) with both phase and momentum relaxation (b) with phase relaxation. Parameters are $E_{F}=10\mu eV$, $E_{M}=0.3\mu eV$, $\Gamma=1\mu eV$, $\phi=\phi_{0}$ and elastic scattering at junction $J_{1},J_{2}$, where $\phi=hc/e$ and $\alpha=\frac{4}{9}$ and represents elastic scattering at J1 and J2.
In Fig. 2, we plot the Lorenz ratio vs. the strength of inelastic scattering for the setup with both phase and momentum relaxation [18], and the setup with only phase relaxation [19] in the presence of coupled MBS, individual MBS, and the absence of MBS. We use the parameters, Fermi energy $E_{F}=10\mu eV$, and flux $\phi=\phi_{0}$, $\phi_{0}$ being the flux quantum $hc/e$. In the absence of MBS, WF law is preserved in all three cases, i.e., $W=1$ for the case of both phase and momentum relaxation (Fig. 2 (a)), and for case of only phase relaxation (Fig. 2 (b)) regardless of inelastic scattering $\epsilon$. For the presence of MBS, we consider two cases: individual MBS and coupled MBS. In our setup, we consider the MBS coupling strength $E_{M}=10\mu eV$ when considering coupled MBS. The violation in Wiedemann-Franz law is highest when $E_{F}$ is close to $E_{M}$. Thus, to distinguish between individual MBS ($E_{M}=0$), and coupled MBS ($E_{M}=3\mu eV$), we take $E_{F}=10\mu eV$.
For $\epsilon\to 0$, the BVT probe is completely disconnected, and there is no inelastic scattering in the setup for both cases. When MBS are uncoupled or individual ($E_{M}=0$), the Lorenz ratio ($W$) is one at the limit of no inelastic scattering and decays with increasing strength of inelastic scattering ($\epsilon$). For the Luttinger liquid setup with individual MBS studied in Ref. [16], the Lorenz ratio $W$ passes through one at $g=1$, i.e., in the absence of inelastic scattering. Thus, our results for individual MBS corroborate the findings of the Luttinger liquid model studied in Ref. [16]. For case with both phase and momentum relaxation [18], the Lorenz ratio scales with inelastic scattering ($\epsilon$) as $W=\frac{0.76}{\epsilon^{1/4}}$. Finally, for case with phase relaxation [19], the Lorenz ratio scales with $\epsilon$ as $W=\frac{0.71}{\epsilon^{1/4}}$. Thus, for individual Majorana, the Lorenz ratio $W$ is inversely proportional to the strength of inelastic scattering $\epsilon$ for inelastic scattering with both phase and momentum relaxation with factor 1/4. Similarly, the Lorenz ratio scales inversely with $\epsilon^{1/4}$ for inelastic scattering with phase relaxation only.
Figure 4: Effect of Majorana bound state energy: Lorentz ratio ($W$) vs. the coupling strength ($\epsilon$) of the BVT probe for (a) with both phase and momentum relaxation (b) with phase relaxation. Parameters are $E_{F}=10\mu eV$, $E_{M}=2.55\mu eV$, $\Gamma=1\mu eV$, $\phi=\phi_{0}$ and elastic scattering at junction $J_{1},J_{2}$, where $\phi=hc/e$ and $\alpha=\frac{1}{2}$ and represents elastic scattering at J1 and J2.
Setup Coupled MBS Individual MBS
Topological Kondo model [16] N/A $W=\frac{0.66}{g}$
MBS ABI with both phase and momentum relaxation
see Eq. (12) and Ref. [18]
$W\approx\frac{0.74}{\epsilon^{1/5}}$ $W\approx\frac{0.76}{\epsilon^{1/4}}$
MBS ABI with only phase relaxation
see Eq. (24) and Ref. [19]
$W\approx\frac{0.72}{\epsilon^{1/5}}$ $W\approx\frac{0.71}{\epsilon^{1/4}}$
For coupled MBS, $E_{M}=0.3\mu eV\neq 0$, we see that the Lorenz ratio is greater than one in the absence of inelastic scattering for all three cases, i.e. $W>1$ for $\epsilon\to 0$ for case of both phase and momentum relaxation [18], and $\epsilon\to 0$ for the case of only phase relaxation. In the presence of coupled MBS, PHS is broken such that the majority of electrons and holes travel in the same direction. It leads to a reduction in the net charge current and a commensurate increase in the heat current. Thus, in the presence of coupled MBS, WF law is violated even in the limit of zero inelastic scattering. The Lorenz ratio decays with increasing $\epsilon$ universally in the presence of MBS, regardless of whether they are coupled or uncoupled. Thus, in the presence of coupled MBS, inelastic scattering can recover the WF law at a particular $\epsilon$. From Fig. 2, we find that the Lorenz ratio $W$ scaling with respect to inelastic scattering $\epsilon$ also follows the power-law for coupled MBS. For case of both phase and momentum relaxation [18] (Fig. 2 (a)), the Lorenz ratio scales with inelastic scattering ($\epsilon$) as $W=\frac{0.74}{\epsilon^{1/5}}$. For the case of only phase relaxation [19], the Lorenz ratio scales with $\epsilon$ as $W=\frac{0.72}{\epsilon^{1/5}}$. Thus, the Lorenz ratio for coupled MBS is inversely proportional to $\epsilon^{1/5}$ when both phase and momentum relaxation are present. In the presence of inelastic scattering with phase relaxation only, the Lorenz ratio is again inversely proportional to $\epsilon^{1/5}$.
VI Analysis
VI.1 Comparison with the work of Buccheri et.al. [16]
We compare our results with the Luttinger liquid model studied in Ref. [16] in Table I. The setup in Ref. [16] considers a quantum dot junction hosting MBS in the topological Kondo regime using the many-body formalism. The Luttinger parameter $g$ describes electron-electron interaction in the many-body picture with $g=1$ corresponding to no interaction. The authors in Ref. [16] report that weakly coupled MBS ($E_{M}\approx 0$) violate WF law. The authors show that the Lorenz ratio scales as $W=\frac{2}{3g}$, i.e., the Lorenz ratio is inversely proportional to the Luttinger parameter in the presence of uncoupled MBS. In contrast in our setup, the Lorentz ratio is inversely proportional to $\epsilon^{1/4}$ for elastic scattering with both phase and momentum relaxation. Similarly, for only phase relaxation case, Lorentz ratio is inversely proportional to $\epsilon^{1/4}$. Table I shows the power law scaling in our setup, either with phase and momentum relaxation or with only phase relaxation. In either case, one identical distinct power law violations for uncoupled and coupled MBS. This provides another unique method to detect MBS in mesoscopic superconducting junctions.
VI.2 Effect of elastic scattering at 3 way junctions J1 and J2
We look for the effects of elastic scattering on the power law scaling. In our calculation till now, we have considered strength of elastic scattering at the 3 way junctions J1 and J2 to be $\alpha=\frac{1}{2}$, see Eq. (4). We can also study the power law scaling for a different value of elastic scattering $\alpha$ at junctions J1 and J2. We only have elastic scattering, as phase of electron is preserved in the scattering and terminals 1 and 2 are current probes. In Fig. 3, we consider $\alpha=\frac{4}{9}$ and we observe that the power law scaling does not change. The reason for using this S-matrix as in Eq. (4) is that it corresponds exactly to a waveguide result, see Ref. [43]. We observe that for the uncoupled MBS case, the Lorentz ratio with both phase and momentum relaxation follows a power law $\frac{0.81}{\epsilon^{1/4}}$, while only for phase relaxation, it follows a power law $\frac{0.80}{\epsilon^{1/4}}$, see Fig. 3. Similarly, for the coupled MBS case, the Lorentz ratio for both phase and momentum relaxation followed a power law of $\frac{0.89}{\epsilon^{1/5}}$, where as for phase relaxation, it is $\frac{0.91}{\epsilon^{1/5}}$, see Fig. 3. We observe that the presence of elastic scattering does affect the prefactor to some extent but the power law in the violation of the Wiedemann-Franz law remains same.
VI.3 Effect of Majorana bound state energy
In Fig. 2, we observe that the coupled MBS has weaker violation compared to the uncoupled MBS. However, this is not always the case. In Fig. 2, we had considered the Majorana bound state energy $E_{M}=0.3\mu eV$ for the coupled MBS. However, if we consider $E_{M}=2.55\mu eV$, then the coupled MBS can have stronger violation than uncoupled MBS ($E_{M}=0$), see, Fig. 4. We observe that while the Lorentz ratio for the phase and momentum relaxation follows a power law of $\frac{0.76}{\epsilon^{1/4}}$, while only for phase relaxation follows a power law of $\frac{0.71}{\epsilon^{1/4}}$ for the uncoupled MBS. For coupled MBS, the Lorentz ratio with phase and momentum relaxation follows a power law $\frac{1.05}{\epsilon^{1/5}}$, where as for only phase relaxation follows a power law $\frac{1.11}{\epsilon^{1/5}}$. Herein also although the power law remains unchanged, the numerical prefactor undergoes a significant change due to change in $E_{M}$ value.
VII Conclusion
We studied the violation of Wiedemann-Franz law in the presence and absence of MBS and the scaling of the Lorenz ratio concerning the strength of inelastic scattering induced by a BVT probe. We find that WF law is only violated in the presence of MBS. We studied the scaling of the Lorenz ratio in the presence of inelastic scattering with both phase and momentum relaxation and with only phase relaxation and compared our results with those of the Luttinger liquid model studied in Ref. [16]. We showed that the Lorenz ratio decays with increasing strength of inelastic scattering when MBS are present, irrespective of whether they are individual or coupled. The Luttinger liquid model predicts that the Lorenz ratio is inversely proportional to the Luttinger parameter with $W(g)=\frac{2}{3g}$. In the presence of individual MBS with both phase and momentum relaxation, the Lorenz ratio scales with inelastic scattering ($\epsilon$) as $W=\frac{0.76}{\epsilon^{1/4}}$, whereas for phase relaxation only, the Lorenz ratio scales with $\epsilon$ as $W=\frac{0.71}{\epsilon^{1/4}}$. When MBS are coupled, the scaling is similar with $W=\frac{0.74}{\epsilon^{1/5}}$ for phase and momentum relaxation, and $W=\frac{0.72}{\epsilon^{1/5}}$ with phase relaxation only [19]. We also observe that the presence of elastic scattering at J1 and J2 and Majorana bound states energy does not have any impact on the power law, but the numerical prefactor significantly changes. The results show that the electron-electron interaction in the Luttinger liquid model is similar to the inelastic scattering induced phenomenologically by the BVT probe [20] with both phase and momentum relaxation [18, 23]. Further, we show that for coupled MBS, the Lorenz ratio is greater than one when inelastic scattering is absent and decays with increasing $\epsilon$. For individual MBS, the Lorenz ratio is conserved without inelastic scattering and decays when inelastic scattering is introduced. This distinct behavior of the Lorenz ratio in the scaling can be used to detect MBS.
There can be several ways of implementing this work experimentally. Refs. [44, 45] show the experimental realization of proximity-induced superconductivity in topological insulators, which is a key component of our setup. The superconductor-ferromagnet junction has also been experimentally realized before [46]. On the question on effect of magnetic field. The magnetic field is shielded such that it does not influence the magnetization of the ferromagnet. This is similar to the Aharonov-Bohm effect seen in normal metal rings and semiconductors. As such, the magnetization in the ferromagnets will be finite. Our setup may be realized in a HgTe quantum well, see [30]. The couplers in our study can represent either elastic scattering at J1, J2 or inelastic scattering at the BVT probe. In our approach, the use of a BVT probe plays a critical role. Experimentally, significant progress has already been made in deriving such BVT probes with the realization of small Ohmic contacts that function as voltage probes, as demonstrated in a pioneering study [34]. A voltage probe is a device through which no net charge current flows, achieved by applying a specific voltage bias. However, our proposal requires a BVT probe, which is slightly more advanced, wherein both the net charge and heat currents are zero. Similarly, the couplers J1 and J2 allow elastic scattering through them. The net charge and heat currents via these couplers are non-zero, in contrast to the other coupler through which the net charge and heat currents are zero. These couplers can also be modeled as a three-way junction, similar to the coupler used for inelastic scattering. These advancements in experimental realizations underscore the feasibility and importance of the BVT probe in understanding fundamental quantum transport phenomena in mesoscopic systems, which are important for considering many-body effects in mesoscopic samples phenomenologically [20, 21, 22].
Appendix A Expression for charge and heat current and Onsager coefficients with phase and momentum relaxation
@@
Using Landauer-Buttiker scattering theory, we can find the charge and heat currents in each terminal of the setup described in Fig. 1 (a). In our work, the left and right terminals are at voltages $V_{1}=0$, and $V_{2}=V$, respectively. Similarly, the temperatures of the left and right terminals are $T_{1}$ and $T_{1}+\Delta T_{R}$ respectively. The BVT probe is at temperature $T+\Delta T$, and voltage $V_{3}$. From Eq. (11), we can write the charge and heat currents for our setup in terms of the Onsager elements as,
@@
We are solving Eqs. (3-8) and using Eqs. (12, 32) for the S-matrix of the BVT probe, one can find the transmission probabilities $\mathcal{T}^{ss;kk}_{ij}$ and $\mathcal{T}^{rs;lk}_{ij}$ in the setup. Plugging the transmission probabilities in Eqs. (32-35), along with Eq. (9), allows us to calculate the charge and heat currents in each terminal. We represent the charge and heat currents in terms of the voltage and temperature biases in a matrix notation below: The charge currents in each terminal due to spin-up electrons and holes are given by,
@@
where $G(X)$ is the electrical conductance for particles with transmission probability $X$, $S(X)$ is the Seebeck coefficient for particles with transmission probability $X$, and $L(X)$ is the thermal conductance for particles with transmission probability $X$. We can calculate the charge and heat currents for phase and momentum relaxation given in Eq. (12) by using the respective S-matrix for the BVT probe [20] when solving for the transmission probabilities $\mathcal{T}^{ss;kk}_{ij}$ and $\mathcal{T}^{rs;lk}_{ij}$. In Eqs. (36-39) the electrical conductance $G(\mathcal{T}^{s;k}_{ij}(E))$ is given by,
@@
Appendix B Expression for charge and heat current and Onsager coefficients with phase relaxation